Determine the pH and pOH of 0.25 L of a solution that is 0.0181 M  boric acid and 0.0391 M sodium borate; pKa for B(OH)3 = 9.0 at 25°C.




1
Expert's answer
2021-09-30T03:25:39-0400

pH=pKa+log"\frac{A-}{HA}"

pH=9.00+log"\frac{0.04}{0.0215}" ]

pH = 9.27

pOH=14-pH

pOH=14-9.27=4.73